OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!!!!!!! What is the value of time, t, for the given values of distance, d, and rate, r? d = 198 km, r = 36 km/h A. 7128 km B. 7128 h C. 5.5 km D. 5.5 h

OpenStudy (anonymous):

@Diamond_Swag11

OpenStudy (anonymous):

Whats up?

OpenStudy (anonymous):

oh hold on

OpenStudy (anonymous):

ok, so first you have to set up the equation.

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

the simple intrest "i" is equal to d*r*t

OpenStudy (anonymous):

@Diamond_Swag11 you aren't supposed to just give them the answer, it's in the code of conduct

OpenStudy (anonymous):

C:

OpenStudy (anonymous):

*facepalm* nvm

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Oh wait hold up

OpenStudy (anonymous):

r=d/t 36=198/t (198)36=t t=7128

OpenStudy (anonymous):

So its 7128h

OpenStudy (anonymous):

r t = d, t = d/r \[\frac{198}{36}=\frac{11 \text{hr}}{2} \]

OpenStudy (anonymous):

so B ?

OpenStudy (anonymous):

Correct! xD

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Probably @robtobey is right

OpenStudy (anonymous):

wait so its d

OpenStudy (anonymous):

yes it is D, now you can close this question because it's answered.

OpenStudy (anonymous):

oops sorry